OpenStudy (anonymous):

HEEEELLLPPPP MEEEE PLLLZZZ!!!! MEDAL FOR WHO EVER EXPLAINS IT BEST PLEASE I DONT UNDERSTAND...

OpenStudy (anonymous):

\[If \frac{X+3 }{ 3 }=\frac{ Y+2 }{ 2 },then \frac{ X }{ 3 }=\]

OpenStudy (dumbcow):

sounds like an emergency ...haha

OpenStudy (dumbcow):

i swear ive helped with this exact same question

OpenStudy (anonymous):

it kinda is the test is coming up soon and the teacher said questions like this are on it...so she gave me a few question to work on and this one is one i dont understand :(

OpenStudy (dumbcow):

hint: split the fractions up example: \[\frac{1+2}{3} = \frac{1}{3} + \frac{2}{3}\]

OpenStudy (anonymous):

how do i split up these thought they are different and i have to find wat x/3 equals

OpenStudy (anonymous):

???

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

really i just showed an example of how to split up fractions with single denominator

OpenStudy (dumbcow):

when you add fractions with same denominator, you simply add the numerators and keep denominator well to split up fraction ... create 2 fractions with same denominator

OpenStudy (anonymous):

im not suppose to have this course i am really bad at geometry or any other math

OpenStudy (dumbcow):

?? this is simple fractions

OpenStudy (anonymous):

\[\frac{ x }{ 3 }, \frac{ 3 }{ 3 }= \frac{ y }{ 2 },\frac{ 2 }{ 2 }\]

OpenStudy (anonymous):

like that

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

does it equal y/2?

OpenStudy (dumbcow):

yes good , but you should be adding them No comma!

OpenStudy (anonymous):

wait you add them or cross multiply? i cross multiplied

OpenStudy (dumbcow):

yes because 3/3 = 1 and 2/2 = 1 subtract 1 from both sides leaving x/3 = y/2

OpenStudy (anonymous):

i cross multiplied and got the same thing:)

OpenStudy (dumbcow):

you could cross-multiply but not in this case because you are trying to isolate " x/3"

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i did x/3 x 3/3 = y/2 x 2/2 3x=9 = 2y=4 3x/3=9/3 = 2y/2=4/2 x=3 = y=2 this is wat i got @dumbcow

OpenStudy (dumbcow):

woah you split this into 2 different equations ..... and i think you don't understand what "cross-multiply" means they are not asking what does " x=" and "y=" question is x/3 = ? you multiplied instead of added ?? \[\frac{x+3}{3} = \frac{y+2}{2}\] \[\frac{x}{3} +\frac{3}{3} = \frac{y}{2}+\frac{2}{2}\] \[\frac{x}{3} +1 = \frac{y}{2} +1\] \[\frac{x}{3} = \frac{y}{2}\]

OpenStudy (anonymous):

oh ok i got it got it yea i see wat i did wrong

OpenStudy (anonymous):

thank you so much :)